OpenStudy (anonymous):

‎17, 25, 32, 37, 47, ?, ? find the next terms.

OpenStudy (anonymous):

62, 82 ??

OpenStudy (anonymous):

explain plz?

OpenStudy (anonymous):

is that the answer u have ?

OpenStudy (anonymous):

no i dont have answer.so am asking...u.

OpenStudy (anonymous):

or it could be.. 67, 117...no logic just guessing...sry

OpenStudy (anonymous):

next terms are 58, 71

OpenStudy (anonymous):

would u plz explain?

OpenStudy (anonymous):

each term has been got by adding up the sum of two digits in the previous term

OpenStudy (anonymous):

17 +8=25 25+7= 32

OpenStudy (anonymous):

wow thats gr8

OpenStudy (anonymous):

hey 17 --> 1+7=8 so 17+8=25 25 --> 2+5=7 so 25+7=32 ...... 47 --> 4+7=11 so 47+11=58